OpenStudy (anonymous):

Determine a function f such that h(x)=f(g(x)), where g(x)=x+2 and h(x)=(x+2)^2. f(x)=_____________. Please explain? Thank youuu :)

OpenStudy (agent0smith):

What happens to go from (x+2) to (x+2)^2?

OpenStudy (anonymous):

it gets squared?

OpenStudy (agent0smith):

So what does your f(x) need to look like?

OpenStudy (agent0smith):

If g(x) = x+2 f(g(x)) = (x+2)^2 then f(x) = ...?

OpenStudy (anonymous):

f(x)=(x+2)^3 ? :/

OpenStudy (agent0smith):

What if f(x) = x^2 f(x+2) = ...?

OpenStudy (anonymous):

x^2 + 2 ?

OpenStudy (agent0smith):

No... you just did a bunch of these, I'm sure you can find that.

OpenStudy (anonymous):

okay... umm f(x)=x^2 f(x+2)= (x+2)^2 ?

OpenStudy (anonymous):

wait so is that the final answer? :o

OpenStudy (agent0smith):

Well what does f(x) have to be?

OpenStudy (agent0smith):

h(x)=f(g(x)), where g(x)=x+2 and h(x)=(x+2)^2. f(x) = ...?

OpenStudy (anonymous):

f(x) has to be x^2 ?

OpenStudy (agent0smith):

Yes.

OpenStudy (anonymous):

thanks again!! :)